# xApp Conflict Mitigation with Scheduler

### v2

@@ 3 SYSTEM MODEL
This study considers an O-RAN cellular network architecture with $B$ O-RUs concurrently serving $U$ users, managed through the RICs within the O-RAN framework. In this setting, a minimum of $n\geq 2$ distinct xApps, denoted as $\{\mathcal{X}_{1},\mathcal{X}_{2},\ldots,\mathcal{X}_{n}\}$, are deployed within the Near-RT RIC to control KPMs in a downlink OFDM scenario. In this model, it is assumed that each O-RU has $R$ available resource block groups (RBGs), bundles of consecutive resource blocks as defined in [43], representing the smallest time-frequency resource to allocate users.
@@
At any time step $t\in\{1,2,\dots,\tilde{t}\}$ within episode $e$, $\delta_{t,e}^{b,r,u}$ is a binary variable that indicates whether O-RU $b\in B$ allocates RBG $r\in R$ to user $u\in U$, where $\delta_{t,e}^{b,r,u}\in\{0,1\},\,\forall b,r,u$ and $\sum_{u\in U}\delta_{t,e}^{b,r,u}=1,\,\forall b,r$. The instantaneous traffic arrivals for each RBG $r$ allocated to user $u$ at time step $t$ in episode $e$, denoted by $\varrho_{t,e}^{b,r,u}$, are assumed to follow a Poisson distribution with mean data arrival rate $d_{e}$, which remains constant throughout the episode.
@@
and $K$ represents the total number of quantization levels, including the 00 level to indicate no transmission on the allocated RBG. Additionally, the system incorporates user mobility, where each user $u$ moves at a constant random speed $v_{e}^{u}\in(V_{\text{min}},V_{\text{max}})$ with a defined $\mathbb{E}[v_{e}^{u}]$ throughout episode $e$, and has a specified probability $\rho$ of altering direction $\Theta_{t,e}^{u}$ in each time slot $t$. At any given time slot $t$ in episode $e$, the position of user $u$, denoted by $(x_{t,e}^{u},y_{t,e}^{u})$, is updated as:
@@
where $\Delta\Theta\in[0,2\pi]$ denotes the change in direction and initial direction of each user at the start of episode $e$, $\Theta_{0,e}^{u}$, is drawn from a uniform distribution over the interval $[0,2\pi]$ (i.e. $\Theta_{0,e}^{u}\sim\mathcal{U}(0,2\pi)$. This ensures that all possible initial headings are equally likely.
The logarithmic normalized channel state information (CSI) of user $u$ for allocated RGB $r$ from O-RU $b$, denoted by $\zeta_{t,e}^{b,r,u}$, is defined as
@@
$\xi_{t,e}^{b,r,u}=\frac{\delta_{t,e}^{b,r,u}\,h_{t,e}^{b,u}\,p_{t,e}^{b,r}}{\sum_{b^{\prime}\in B,b^{\prime}\neq b}\sum_{u^{\prime}\in U^{b^{\prime}}}\delta_{t,e}^{b^{\prime},r,u^{\prime}}\,h_{t,e}^{b^{\prime},u}\,p_{t,e}^{b^{\prime},r}+\sigma^{2}},$ (6)
@@
$C_{t,e}^{b,r,u}=\mathcal{B}^{r}\log_{2}\left(1+\xi_{t,e}^{b,r,u}\right),$ (7)
where $\mathcal{B}^{r}$ denotes the bandwidth of RBG $r$. The transmission rate $\Psi_{t,e}^{b,r,u}$ equals to $C_{t,e}^{b,r,u}$ if the length of the queued data exceeds the transmission capacity within a time slot, in which case any data beyond that capacity is assumed to be discarded. Otherwise, (i.e. $C_{t,e}^{b,r,u}T\geq\sum_{u\in U}\delta_{t,e}^{b,r,u}\varrho_{t,e}^{b,r,u}$), and the transmission rate is given by $\Psi_{t,e}^{b,r,u}=\varrho_{t,e}^{b,r,u}/T$, where $T$ represents the duration of the time slot $t$.
3.1 From REINFORCE to Advantage Actor-Critic (A2C)
@@
where $\pi_{\theta}(\mathbf{a}_{t}\mid\mathbf{s}_{t})$ is the probability of taking action $\mathbf{a}_{t}$ in state $\mathbf{s}_{t}$ and and $G_{t}=\sum_{j=t}^{\tilde{t}-1}\gamma^{j-t}\tau_{j}$ denotes the return with $\gamma$ is the discount factor and reward $\tau$ following time step $t$. This leads to update the policy parameters $\theta$ via
$\theta\leftarrow\theta+\eta\sum_{t=0}^{\tilde{t}-1}\gamma^{t}G_{t}\,\nabla_{\theta}\log\pi_{\theta}(\mathbf{a}_{t}\mid\mathbf{s}_{t}),$ (9)
where $\eta$ is learning rate. While this Monte Carlo approach is conceptually straightforward, it exhibits high variance and necessitates waiting until an episode terminates [47]. A common variance reduction technique is to subtract a state-dependent baseline $b(s_{t})$ [48], which does not depend on the chosen action $a_{t}$, leading to the unbiased gradient
@@
Replacing the full return $G_{t}$ with the one-step return $G_{t:t+1}=r_{t}+\gamma V_{\phi}(\mathbf{s}_{t+1})$ further decreases variance by bootstrapping from the critic’s estimate. This yields the advantage,
$A_{t}=r_{t}+\gamma V_{\phi}(\mathbf{s}_{t+1})\;-\;V_{\phi}(\mathbf{s}_{t}),$ (12)
@@
$\theta\leftarrow\theta+\eta\sum_{t=0}^{\tilde{t}-1}\gamma^{t}\,\bigl{[}G_{t:t+1}-V_{\phi}(\mathbf{s}_{t})\bigr{]}\,\nabla_{\theta}\log\pi_{\theta}(\mathbf{a}_{t}\mid\mathbf{s}_{t}).$ (14)
In A2C approach, a synchronous variant of the Asynchronous Advantage Actor-Critic (A3C) algorithm introduced in [51], a policy network (the Actor), $\pi_{\theta}(\mathbf{a}_{t}\mid\mathbf{s}_{t})$, outputs a probability distribution over actions $\mathbf{a}_{t}$ given state $\mathbf{s}_{t}$ while a value network (the Critic), $V_{\phi}$, estimates the expected return from state $\mathbf{s}_{t}$, $V_{\phi}(\mathbf{s}_{t})\approx\mathbb{E}[\,G_{t}\mid\mathbf{s}_{t}\,]$. The advantage function, $A(\mathbf{s}_{t},\mathbf{a}_{t})=G_{t:t+1}-V_{\phi}(\mathbf{s}_{t})$, compares the discounted one-step return $G_{t:t+1}$ to the Critic’s baseline $V_{\phi}(\mathbf{s}_{t})$. By quantifying how much better (or worse) a particular action is relative to the baseline, this formulation reduces the variance of policy gradient updates while maintaining low bias, thereby enhancing learning stability. The policy gradient update to $\theta$ aims to maximize the expected advantage
$\sum_{t=0}^{\tilde{t}-1}A_{t}\;\log\pi_{\theta}\bigl{(}\mathbf{a}_{t}\mid\mathbf{s}_{t}\bigr{)},$ (15)
@@
$\sum_{t=0}^{\tilde{t}-1}\Bigl{(}G_{t}-V_{\phi}(\mathbf{s}_{t})\Bigr{)}^{2},$ (16)
leading to a combined loss function:
$\mathcal{L}(\theta,\phi)\;=\;-\,\sum_{t=0}^{\tilde{t}-1}\Bigl{[}A_{t}\,\log\pi_{\theta}\bigl{(}\mathbf{a}_{t}\mid\mathbf{s}_{t}\bigr{)}\Bigr{]}\;+\;\alpha\,\sum_{t=0}^{\tilde{t}-1}\Bigl{(}G_{t}-V_{\phi}(\mathbf{s}_{t})\Bigr{)}^{2},$ (17)
@@
3.2 Power and RGB Allocation xApps with A2C
@@
where $\mathbf{s}_{t,e}$ contains four features for every RGBs of each O-RU assigned to the users. The action vector, representing the power allocations for all RBGs, is given by:
@@
In A2C approach, a policy network (the Actor), $\pi_{\theta}$, outputs a probability distribution over possible power levels for each RGB per O-RU, while a value network (the Critic), $V_{\phi}$, estimates the expected return for a given state. The joint policy of independent action selection per RBG, $a^{r}$,
@@
$G_{t,e}=\sum_{j=0}^{\tilde{t}-t-1}\gamma^{j}\,\tau_{t+j,e},\quad\gamma\in(0,1).$ (22)
Here, $\tilde{t}$ is the final timestep within an episode, $e$, and $\tau_{t+j}$ is the reward at time $t+j$.
Resource allocation in O-RAN has been examined from the perspective of physical resource blocks using a variety of methods; random forest classifier [58], probabilistic forecasting techniques [59], DRL [60], A2C [61], DQN [24]. In this work, the DQN-based radio resource allocation xApp presented in [40] is re-designed as an RGB allocation xApp, wherein the A2C method is employed to assign RBGs to users across O-RUs. This procedure parallels that of the power allocation xApp, wherein an Actor network outputs a probability distribution over RBG assignment decisions
@@
4 SCHEDULER DRIVEN xApp MANAGEMENT FRAMEWORK
@@ 4.3 Scheduler
Upon receiving an intent from the MNO, the Non-RT RIC processes and selects appropriate xApps, $\{\mathcal{X}_{1},\mathcal{X}_{2},\ldots,\mathcal{X}_{n}\}$, according to the target objective of intent, $f$, (e.g., maximizing data rate) and provides $\varepsilon$ context variables, $c_{\dagger}=\left\{c_{\dagger}^{1},c_{\dagger}^{2},\dots,c_{\dagger}^{\varepsilon}\right\}$ as EI via A1 for each scheduling period $\dagger$. The scheduler is also modelled using the A2C algorithm, incorporating context variables $c_{\dagger}$ and reward $f_{\dagger}$ as inputs to determine the active xApps for the subsequent scheduling period. The scheduler sends an activation message, $\left\{\mu_{\dagger}^{1},\mu_{\dagger}^{2},\dots,\mu_{\dagger}^{n}\mid\mu_{\dagger}^{n}\in\{0,1\}\right\}$, is to xApps IH, thereby enabling the selected xApps to access the data repository and commence optimization based on their respective objectives. In other words, the Actor network for scheduler A2C model, $\pi_{\theta}(\mathbf{a}_{\dagger}\mid\mathbf{s}_{\dagger})$, outputs a probability distribution over activation decisions, where the state vector $\mathbf{s}_{\dagger}=\left[c_{\dagger}^{1},\ c_{\dagger}^{2},\ \dots,\ c_{\dagger}^{\varepsilon},\ f_{\dagger}\right]$ and the action vector $\mathbf{a}_{\dagger}=\left[\mu_{\dagger}^{1},\ \mu_{\dagger}^{2},\ \dots,\ \mu_{\dagger}^{n}\right]$.
5 PERFORMANCE EVALUATION
@@
Table 1 presents the parameters of the simulation environment, which was created in Python to facilitate the training and testing of xApps and the scheduler, and which employs an O-RAN mobile network with four O-RUs concurrently serving 16 users.
5.1 Training of Power and RGB Allocation xApps
Training of the Power and RGB xApps is conducted in episodes, and the simulation environment is reset at the beginning of each episode ($t=0$); the users are evenly distributed among four O-RUs, and each user $u\in U$ is randomly positioned at a distance between $150$ and $450$ m from their assigned O-RU. Both xApps are trained under the assumption of equal allocation of all other resources, which serves as a baseline allocation strategy. For instance, the available $12$ RGBs for each O-RU are evenly distributed among connected users during the training of the Power xApp. For each episode $e$, the mean data arrival rate $d_{e}$ and the average user speed, $\frac{1}{U}\sum_{u=1}^{U}v_{e}^{u}$, are randomly selected from the finite sets $\{3,5,7,9\}$ Mbps and $\{10,20,30,40\}$ m/s, respectively. The objective of both xApps is to maximize the normalized data transmission for each episode, $\tau_{e}$, calculated as
$\tau_{e}=\frac{\sum_{t=0}^{\tilde{t}-1}\tau_{t,e}}{d_{e}\times R\times B},\quad\tau_{t,e}=\sum_{b=1}^{B}\sum_{r=1}^{R}\sum_{u=1}^{U}\delta_{t,e}^{b,r,u}\Psi_{t,e}^{b,r,u},$ (25)
where $\tau_{t,e}$ is the total data rate transmitted across all RBGs ($R\times B$) during each time step $t$ in episode $e$ and the maximum achievable data transmission is defined as the product of the average load $d_{e}$ and the total number of RBGs. The training results of the Power and RGB xApps are presented in Fig. 4, evaluated over $100,000$ episodes, with each episode $e$ comprising $\tilde{t}=50$ time steps.
(a) Training of Power xApp
(b) Training of RGB xApp
@@ 5.1 Training of Power and RGB Allocation xApps
Figure 4: Moving average reward of both xApps with a sliding window of 500
@@
Fig. 5 illustrates the performance of the trained xApps—specifically, RBG-only, Power-only, and Both (i.e., when deployed independently)—across various scenarios defined by mean data arrival rates (low: $2$ Mbps, mid: $5$ Mbps, high: $8$ Mbps) and average user speeds (low: $5$ m/s, mid: $25$ m/s, high: $45$ m/s). Two primary observations can be drawn from the analysis. First, the independent control of parameters by concurrently deployed xApps, in the absence of coordination, leads to performance degradation across all scenarios. Second, the severity of these conflicts depends on network state variables; notably, performance degradation reaches $16\%$ in the high scenario ($8$ Mbps, $45$ m/s), compared to $5\%$ in the low scenario ($2$ Mbps, $5$ m/s). The conflicting behavior of these xApps can be attributed to their independent resource allocation actions. For example, conflicts may occur when the Power xApp assigns a specific power level to an RBG that is not allocated to any user by the RGB xApp. Conversely, conflicts may also arise when the RGB xApp assigns a substantial number of RBGs to a user, while the Power xApp provides a relatively low power allocation for those RBGs.
Table 1: Simulation Parameters
@@
Min transmission power ($P_{max}$) $1$ dBm
Additive white Gaussian noise $-114$ dBm
Propagation model $120.9+37.6\log(\upsilon)$ dB
Log-Normal shadowing $8$ dB
@@
Time slots per episode ($\tilde{t}$) $50$
@@
5.2 Training of Scheduler
The initial condition for power and RGB allocation at the beginning of each episode is assumed to be an equal division, as established during the individual training of the xApps. Although the scheduler can be configured to optimize multiple objectives, this study specifically focuses on maximizing the normalized transmission rate achieved by both xApps. Additionally, the reward for the scheduler is computed on an episodic basis over =10$\dagger$=10† = 10 time steps.
Method 1: Keep previous action

### v3

@@ III SYSTEM MODEL
This study considers an O-RAN cellular architecture with a set of O-RUs, denoted by $\mathcal{B}=\{1,\dots,B\}$, concurrently serving a user set $\mathcal{U}=\{1,\dots,U\}$, managed through the RICs within the O-RAN framework. In this setting, a minimum of $n\geq 2$ distinct xApps, denoted as $\{\mathcal{X}_{1},\mathcal{X}_{2},\ldots,\mathcal{X}_{n}\}$, are deployed within the Near-RT RIC to control KPMs in a downlink OFDM scenario. In this model, it is assumed that each O-RU possesses a set of available resource block groups (RBGs), $\mathcal{R}=\{1,\dots,R\}$, which are bundles of consecutive resource blocks as defined in [95] and represent the smallest allocable time-frequency resources for user assignments.
@@
At any time step $t\in\{1,2,\dots,T\}$ within episode $e$, $\delta_{t,e}^{b,r,u}$ is a binary variable that indicates whether O-RU $b\in B$ allocates RBG $r\in R$ to user $u\in U$, where $\delta_{t,e}^{b,r,u}\in\{0,1\},\,\forall b,r,u$ and $\sum_{u\in U}\delta_{t,e}^{b,r,u}=1,\,\forall b,r$. The instantaneous traffic arrivals for each RBG $r$ allocated to user $u$ at time step $t$ in episode $e$, denoted by $\varrho_{t,e}^{b,r,u}$, are assumed to follow a Poisson distribution with mean data arrival rate $d_{e}$, which remains constant throughout the episode.
@@
and $|\mathcal{P}|=K$ is the total number of quantization levels, including the 0 level to indicate no transmission on the allocated RBG.
Additionally, the system incorporates user mobility, wherein each user $u$ moves at a constant random speed $v_{e}^{u}\in[V_{\text{min}},V_{\text{max}}]$ throughout episode $e$, characterized by a mean speed $\mathbb{E}[v_{e}^{u}]$, and each user $u$ has a specified probability $\rho$ of altering direction $\Theta_{t,e}^{u}$ in each time slot $t$. At any given time slot $t$ in episode $e$, the position of user $u$, denoted by $(x_{t,e}^{u},y_{t,e}^{u})$, is updated as:
@@
where $\Delta\Theta\in[0,2\pi]$ denotes the change in direction and initial direction of each user at the start of episode $e$, $\Theta_{0,e}^{u}$, is drawn from a uniform distribution over the interval $0,2\pi]$.
The logarithmic normalized channel state information (CSI) of user $u$ for allocated RBG $r$ from O-RU $b$, denoted by $\zeta_{t,e}^{b,r,u}$, is defined as
@@
$\xi_{t,e}^{b,r,u}=\frac{\delta_{t,e}^{b,r,u}\,h_{t,e}^{b,u}\,p_{t,e}^{b,r,u}}{\sum_{b^{\prime}\in B,b^{\prime}\neq b}\sum_{u^{\prime}\in U}\delta_{t,e}^{b^{\prime},r,u^{\prime}}\,h_{t,e}^{b^{\prime},u}\,p_{t,e}^{b^{\prime},r,u}+\sigma^{2}},$ (6)
@@
$C_{t,e}^{b,r,u}=W^{r}\log_{2}\left(1+\xi_{t,e}^{b,r,u}\right),$ (7)
where $W^{r}$ denotes the bandwidth of RBG $r$. The transmission rate $\Psi_{t,e}^{b,r,u}$ equals to $C_{t,e}^{b,r,u}$ if the length of the queued data exceeds the transmission capacity within a time slot, in which case any data beyond that capacity is assumed to be discarded. Otherwise, (i.e. $C_{t,e}^{b,r,u}T_{s}\geq\sum_{u\in U}\delta_{t,e}^{b,r,u}\varrho_{t,e}^{b,r,u}$), and the transmission rate is given by $\Psi_{t,e}^{b,r,u}=\varrho_{t,e}^{b,r,u}/T_{s}$, where $T_{s}$ represents the duration of the time slot $t$.
III-A From REINFORCE to Advantage Actor-Critic (A2C)
@@
where $\pi_{\theta}(\mathbf{a}_{t}\mid\mathbf{s}_{t})$ is the probability of taking action $\mathbf{a}_{t}$ in state $\mathbf{s}_{t}$ and $G_{t}=\sum_{j=t}^{T-1}\gamma^{j-t}\tau_{j}$ denotes the return with $\gamma$ is the discount factor and reward $\tau$ following time step $t$. This leads to update the policy parameters $\theta$ via
$\theta\leftarrow\theta+\eta\sum_{t=0}^{T-1}G_{t}\,\nabla_{\theta}\log\pi_{\theta}(\mathbf{a}_{t}\mid\mathbf{s}_{t}),$ (9)
where $\eta$ is the learning rate. While this Monte Carlo approach is conceptually straightforward, it exhibits high variance and necessitates waiting until an episode terminates [99]. A common variance reduction technique is to subtract a state-dependent baseline $b(\mathbf{s}_{t})$ [100], which does not depend on the chosen action $a_{t}$, leading to the unbiased gradient
@@
Replacing the full return $G_{t}$ with the one-step return $G_{t:t+1}=\tau_{t}+\gamma V_{\phi}(\mathbf{s}_{t+1})$ further decreases variance by bootstrapping from the critic’s estimate. This yields the advantage,
$A_{t}=\tau_{t}+\gamma V_{\phi}(\mathbf{s}_{t+1})\;-\;V_{\phi}(\mathbf{s}_{t}),$ (12)
@@
$\theta\leftarrow\theta+\eta\sum_{t=0}^{T-1}\,\bigl{[}G_{t:t+1}-V_{\phi}(\mathbf{s}_{t})\bigr{]}\,\nabla_{\theta}\log\pi_{\theta}(\mathbf{a}_{t}\mid\mathbf{s}_{t}).$ (14)
In the A2C approach, a synchronous variant of the Asynchronous Advantage Actor-Critic (A3C) algorithm introduced in [103], a policy network (the Actor), $\pi_{\theta}(\mathbf{a}_{t}\mid\mathbf{s}_{t})$, outputs a probability distribution over actions $\mathbf{a}_{t}$ given state $\mathbf{s}_{t}$ while a value network (the Critic), $V_{\phi}$, estimates the expected return from state $\mathbf{s}_{t}$; $V_{\phi}(\mathbf{s}_{t})\approx\mathbb{E}[\,G_{t}\mid\mathbf{s}_{t}\,]$. The advantage function, $A(\mathbf{s}_{t},\mathbf{a}_{t})=G_{t:t+1}-V_{\phi}(\mathbf{s}_{t})$, compares the discounted one-step return $G_{t:t+1}$ to the Critic’s baseline $V_{\phi}(\mathbf{s}_{t})$. By quantifying how much better (or worse) a particular action is relative to the baseline, this formulation reduces the variance of policy gradient updates while maintaining low bias, thereby enhancing learning stability. The policy gradient update to $\theta$ aims to maximize the expected advantage
$\sum_{t=0}^{T-1}A_{t}\;\log\pi_{\theta}\bigl{(}\mathbf{a}_{t}\mid\mathbf{s}_{t}\bigr{)},$ (15)
@@
$\mathcal{L}(\theta,\phi)=-\sum_{t=0}^{T-1}A_{t}\,\log\pi_{\theta}\bigl{(}\mathbf{a}_{t}\mid\mathbf{s}_{t}\bigr{)}+\alpha\sum_{t=0}^{T-1}\bigl{(}G_{t}-V_{\phi}(\mathbf{s}_{t})\bigr{)}^{2}\,$
@@
III-B Power and RBG Allocation xApps with A2C
@@
where $\mathbf{s}_{t,e}$ contains four features for every RBGs of each O-RU assigned to the users. The action vector, representing the power allocations for all RBGs, is given by:
@@
In A2C approach, a policy network (the Actor), $\pi_{\theta}$, outputs a probability distribution over possible power levels for each RBG per O-RU, while a value network (the Critic), $V_{\phi}$, estimates the expected return for a given state. The joint policy of independent action selection per RBG, $a^{r}$,
@@
$G_{t,e}=\sum_{j=0}^{T-t-1}\gamma^{j}\,\tau_{t+j,e},\quad\gamma\in(0,1).$ (21)
Here, $T$ is the final timestep within an episode, $e$, and $\tau_{t+j}$ is the reward at time $t+j$.
Resource allocation in O-RAN has been examined from the perspective of physical resource blocks using a variety of methods; random forest classifier [110], probabilistic forecasting techniques [111], DRL [112], A2C [113], DQN [76]. In this work, the DQN-based radio resource allocation xApp presented in [92] is re-designed as an RBG allocation xApp, wherein the A2C method is employed to assign RBGs to users across O-RUs. This procedure parallels that of the power allocation xApp, wherein an Actor network outputs a probability distribution over RBG assignment decisions
@@
IV SCHEDULER DRIVEN xApp MANAGEMENT FRAMEWORK
@@ IV-C Scheduler
Upon receiving an intent from the MNO, the Non-RT RIC processes and selects appropriate xApps, $\{\mathcal{X}_{1},\mathcal{X}_{2},\ldots,\mathcal{X}_{n}\}$, according to the target objective of intent, $f_{\dagger}$, (e.g., maximizing data rate) and provides $\varepsilon$ context variables, $c_{\dagger}=\left\{c_{\dagger}^{1},c_{\dagger}^{2},\dots,c_{\dagger}^{\varepsilon}\right\}$ as EI via A1 for each scheduling period $\dagger$. The scheduler sends an activation message, $\left\{\mu_{\dagger}^{1},\mu_{\dagger}^{2},\dots,\mu_{\dagger}^{n}\mid\mu_{\dagger}^{n}\in\{0,1\}\right\}$, is to xApps IH, thereby enabling the selected xApps to access the data repository and commence optimization based on their respective objectives.
V PERFORMANCE EVALUATION
@@
Table I presents the parameters of the simulation environment, which was created in Python to facilitate the training and testing of xApps and the scheduler, and which employs an O-RAN mobile network with four O-RUs concurrently serving 16 users.
V-A Training of Power and RBG Allocation xApps
(a) Training performance of Power xApp
(b) Training performance of RBG xApp
@@ V-A Training of Power and RBG Allocation xApps
Figure 4: Moving average (MA) reward of both xApps with a sliding window of 500
Training of the Power and RBG xApps is conducted in episodes, and the simulation environment is reset at the beginning of each episode ($t=0$); the users are evenly distributed among four O-RUs, and each user $u\in U$ is randomly positioned at a distance between $150$ and $450$ m from their assigned O-RU. The propagation environment is modelled by the large-scale fading coefficient of $120.9+37.6\log_{10}(\omega)+10\log_{10}(Z)\;\text{dB},$ where $\omega$ is the O-RU-to-user distance (in km) and $Z$ is a log-normal random variable with a standard deviation of $8\;\text{dB}$. Both xApps are trained under the assumption of equal allocation of all other resources, which serves as a baseline allocation strategy. For instance, the available $12$ RBGs for each O-RU are evenly distributed among connected users during the training of the Power xApp. For each episode $e$, the mean data arrival rate $d_{e}$ and the average user speed $\mathbb{E}[v_{e}^{u}]$ are randomly selected from the finite sets $\{3,5,7,9\}$ Mbps and $\{10,20,30,40\}$ m/s, respectively. The objective of both xApps is to maximize the normalized data transmission for each episode, $\tau_{e}$, calculated as
$\tau_{e}=\frac{\sum_{t=0}^{T-1}\tau_{t,e}}{d_{e}\times R\times B},\quad\tau_{t,e}=\sum_{b=1}^{B}\sum_{r=1}^{R}\sum_{u=1}^{U}\delta_{t,e}^{b,r,u}\Psi_{t,e}^{b,r,u},$ (24)
where $\tau_{t,e}$ is the total data rate transmitted across all RBGs ($R\times B$) during each time step $t$ in episode $e$ and the maximum achievable data transmission is defined as the product of the average load $d_{e}$ and the total number of RBGs. The training results of the Power and RBG xApps are presented in Fig. 4, evaluated over $100,000$ episodes, with each episode $e$ comprising $T=50$ time steps.
@@
TABLE I: Simulation Parameters
@@
Min transmission power ($P_{max}$) 11 dBm
Additive white Gaussian noise 114-114 dBm
Propagation model $120.9+37.6\log(\omega)$ dB
Log-Normal shadowing 88 dB
@@
Time slots per episode ($T$) $50$
@@
Fig. 5 illustrates the performance of the trained xApps—specifically, RBG-only, Power-only, and Both (i.e., when deployed independently)—across various scenarios defined by mean data arrival rates (low: $2$ Mbps, mid: $5$ Mbps, high: $8$ Mbps) and average user speeds (low: $5$ m/s, mid: $25$ m/s, high: $45$ m/s). Two primary observations can be drawn from the analysis. First, the independent control of parameters by concurrently deployed xApps, in the absence of coordination, leads to performance degradation across all scenarios. Second, the severity of these conflicts depends on network state variables (i.e., context variables); notably, performance degradation reaches $16\%$ in the high scenario ($8$ Mbps, $45$ m/s), compared to $5\%$ in the low scenario ($2$ Mbps, $5$ m/s). The conflicting behavior of these xApps can be attributed to their independent resource allocation actions. For example, conflicts may occur when the Power xApp assigns a specific power level to an RBG that is not allocated to any user by the RBG xApp. Conversely, conflicts may also arise when the RBG xApp assigns a substantial number of RBGs to a user, while the Power xApp provides a relatively low power allocation for those RBGs.
V-B Training of Scheduler
The scheduler is also modelled using the A2C algorithm, incorporating context variables $c_{\dagger}$ and reward $f_{\dagger}$ as inputs to determine the active xApps for the subsequent scheduling period. In other words, the Actor network for scheduler A2C model, $\pi_{\theta}(\mathbf{a}_{\dagger}\mid\mathbf{s}_{\dagger})$, outputs a probability distribution over activation decisions, where the state vector $\mathbf{s}_{\dagger}=\left[c_{\dagger}^{1},\ c_{\dagger}^{2},\ \dots,\ c_{\dagger}^{\varepsilon},\ f_{\dagger}\right]$ and the action vector $\mathbf{a}_{\dagger}=\left[\mu_{\dagger}^{1},\ \mu_{\dagger}^{2},\ \dots,\ \mu_{\dagger}^{n}\right]$. The initial condition for power and RBG allocation at the beginning of each episode is assumed to be an equal division, consistent with the settings established during the individual training of the xApps. Furthermore, the reward for the scheduler is computed episodically over $\dagger=10$} time steps. Although the scheduler can be configured to optimize multiple objectives, the current study specifically focuses on maximizing the normalized transmission rate, $\tau_{e}$, achieved by both xApps. Note that only two contextual variables are used; adding more could further improve scheduler performance by capturing a more comprehensive picture of the network’s operational state. This framework eliminates the need for any additional offline re-training or joint training of the xApps themselves; only the lightweight scheduler is re-trained online according to intent and the group of xApps that are employed.
Method 1: Retain previous action
In this method, the training of the scheduler is carried out with two pre-trained A2C xApps— the power ($\mathcal{X}_{1}$) and RBG ($\mathcal{X}_{2}$)—to decide which xApp(s) to activate, choosing between individual or simultaneous deployment of these xApps. The scheduler is designed to make multiple decisions within a single episode (i.e., five decisions per episode (${\dagger}=10$ time steps)) and share these decisions with xApps through activation messages, $\mu_{\dagger}=\{\mu_{\dagger}^{1},\mu_{\dagger}^{2}\}$ where $\mu_{\dagger}^{n}\in\{0,1\}\quad\forall\,n=1,2$. Two context variables, the average user speed ($c_{\dagger}^{1}$) and mean data arrival rate ($c_{\dagger}^{2}$), are utilized as the state elements of the scheduler for each scheduling period $\dagger$} during the training. This iterative approach enables the scheduler to refine its activation strategy continuously. Furthermore, if only one xApp is activated at a given decision interval, the system retains the last action made by the previously deactivated xApp, ensuring operational consistency throughout the training. In other words, the scheduler prioritizes the set of xApp actions to be applied to the network and iteratively refines its activation strategy during training. Therefore, the method is regarded as an enhanced, A2C-driven form of prioritization [87] and is considered state-of-the-art.
Method 2: Extend with baselines